# Intermittency and non-universality of pair dispersion in isothermal compressible turbulence

### v2

@@ 2 Model and its numerical simulation
which is the ratio of the total kinetic energy in the solenoidal modes to the total kinetic energy. All these parameters of our DNS are summarized in Table 1. We use $L$ as our unit of length, and $T_{\rm E}$, the large-eddy turnover time, as our unit of time. Further details about the DNS are given in Appendix A.
S1 1 0.05 0.053 $1.05$ $6.7\times 10^{3}$ 1.66 32 $0.82$ $8.1\times 10^{-6}$ 0.29 306 0.0111
S2 1 0.02 0.050 $2.5$ $6.3\times 10^{3}$ 1.20 24 $0.68$ $9.6\times 10^{-6}$ 0.26 255 0.0107
C1 0 0.05 0.038 $0.75$ $4.7\times 10^{3}$ 1.04 28 $0.19$ $1.3\times 10^{-5}$ 0.16 123 0.0099
C2 0 0.02 0.040 $2.0$ $5.0\times 10^{3}$ 1.17 29 $0.34$ $8.7\times 10^{-6}$ 0.21 172 0.0109
3 Results
Figure 1: Typical snapshots of $\nabla\cdot{\bf u}$, vorticity ${\bm{\omega}}=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, in the cases (a) S1, (b) S2, (c) C1 and (d) C2. In each plot, ${\bm{\nabla}}\cdot\bm{u}$ and ${\bm{\omega}}$ are normalized by their respective root-mean-square (rms) values. Shocks are clearly visible in the ${\bm{\nabla}}\cdot\bm{u}$ profiles as the blue filament–like structures. The vorticity field, $\omega$, is smoother in S1 than in S2 which contains nearly one-dimensional structures of intense vorticity bordering the shocks; in all figures, the areas of high vorticity lie in the vicinity of shocks. Similarly the spatial distribution of $\omega$ is more homogeneous in C1 than in C2. The values of $\rho$ are clearly enhanced near the shocks; for a given type of external forcing, $\rho$ tends to take much higher values at higher Mach numbers, Ma.
We perform four sets of DNSs, two for $\xi=1$, S1 and S2, and two for $\xi=0$, C1 and C2. Typically, subsonic and supersonic compressible flow can be qualitatively different. Hence we perform a set of simulations that are transonic, Runs S1 and C1, $\mbox{Ma}\simeq 1$, and another set S2 and C2 that are supersonic, $\mbox{Ma}>1$. In Fig. 1 we show pseudocolour plots of ${\bm{\nabla}}\cdot\bm{u}$, ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$ for all the cases or runs. The shocks are clearly visible as filamentary structures with large negative values of ${\bm{\nabla}}\cdot\bm{u}$. Naively, we may expect that the presence of the shocks would not show up in pseudocolour plot of ${\bm{\omega}}$, but this not the case – the shocks can be seen as almost one-dimensional manifolds with large absolute values of the vorticity: This is because, although a shock front that is a straight line does not contribute to vorticity, curved shock fronts do. In our simulations, (a) the density fluctuations are larger for the supersonic cases compared to the transonic ones and (b) the vorticity is more uniformly distributed in the transonic cases than the supersonic ones, for a given type of external forcing, as observed in Fig. 1.
@@ 3 Results
The scaling relations hold for $\eta\ll r\ll L_{\rm I}$, where $\eta$ is the dissipation scale and $L_{\rm I}$ is the integral scale. All the exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ (Konstandin et al., 2012; Wang et al., 2017, 2012b) (see Appendix B.2).
3.1 Pair dispersion
@@
In order to study pair dispersion, we investigate how the separation, $R(t)$, between a pair of nearest-neighbor tracers, separated initially by the grid-spacing (approximately $0.3\eta$), evolves with time. We calculate $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations. Our results are shown in Fig. (2a). At early times, $\left\langle R^{2}(t)\right\rangle$ grows exponentially for all the runs [For S1 and S2 we show the plots in the inset of Fig. (2a)]. This range of exponential growth is very small for C1 and C2. Similar results (Jullien, 2003) have been obtained for incompressible 2D turbulence. Exponential growth up to a length scale typically implies that the flow across the tracers is smooth up to that scale. This holds for our simulations because at such small time scales most of the pair separations lie away from shocks. We note that at late times the behavior of $\left\langle R^{2}\right\rangle$ is not universal, it depends on precisely how the DNS is forced.

### v3

@@ B.1 Parameters and Snapshots
which is the ratio of the total kinetic energy in the solenoidal modes to the total kinetic energy.
In two-dimensional incompressible turbulence, forcing at an intermediate scale gives rise to two cascades: (a) an inverse cascade to large scales and (b) a direct cascade to small scales. The statistics in the inverse cascade is close to Gaussian, whereas the statistics in the direct cascade is intermittent. In two dimensional compressible turbulence, forced at intermediate scales, kinetic energy shows a dual cascade [35]. We choose to force at the large scales, $k_{\rm f}=3$, such that we obtain a large range of direct cascade, see Fig. (4). The statistics of velocity fluctuations in this range is intermittent, as we show later by evaluating the scaling exponents of the $p$-th order structure functions.
We display the three shell-averaged velocity spectra,
$E(k)$ $=|\hat{{\bm{u}}}({\bm{k}})|^{2}\/,$ (14a)
$E_{\rm s}(k)$ $=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}\/,$ (14b)
$\text{and}\quad E_{\rm c}(k)$ $=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$ (14c)
respective for the total, solenoidal, and compressive velocity. They are plotted in Fig. (5). In all the runs, $E_{\rm c}\sim k^{-2}$ over a significant range of scales. Earlier works [35] have also observed this over a smaller range of scales and have identified it with acoustic turbulence. In S1, $E_{\rm s}(k)$ falls off much faster than that in S2. Thus, the small-scale fluctuations of $\bm{u}^{\rm s}$ are much smoother in S1 than those in S2, and this is a reflection of the elongated intense vorticity patches, which are aplenty in S2, but not in S1. In S2, $E_{\rm s}(k)$ appears to scale as $k^{-3}$, across a small regime of $k>k_{\rm f}$. In the irrotationally-forced runs, $E_{\rm c}(k)\gg E_{s}(k)$ across whole of the inertial range with the difference being small at higher Ma. $E_{\rm s}(k)$, however, does not appear to have a well-defined power-law regime.
B.3 Equal-time structure functions
Figure 6: Equal-time structure functions. Log-log plots of $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, defined in  (15) versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2. The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$. In each figure, the data has been averaged over a single non-equilibrium steady state (NESS) snapshot.
Figure 7: Equal-time exponents $\zeta_{p}$ for the four runs.
@@ B.3 Equal-time structure functions
The scaling relations hold for $\eta\ll r\ll L_{\rm I}$, where $\eta$ is the dissipation scale and $L_{\rm I}$ is the integral scale. All the exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ [29, 31, 32].
We show plots of various equal-time longitudinal structure functions of the total velocity, $S_{\rm p}(r)$, and its solenoidal, $S^{\rm s}_{\rm p}(r)$, and irrotational, $S^{\rm c}_{\rm p}(r)$, components in Fig. (6). In each image, the regime of power-law fitting has been demarcated by dashed lines. In Figs. 6(a) and 6(b), note that the scaling regime of $S^{\rm s}_{\rm p}(r)$ in S2 occurs across higher values of $r$ than that in S1. Moreover, for a given $p$, the slope in the scaling-range of $S^{\rm s}_{\rm p}(r)$ in the former is visibly smaller than that of $S^{\rm s}_{\rm p}(r)$ in the latter. For $S_{\rm 2}^{\rm s}(r)$, the nature of this change can also be inferred from the scaling of $E_{\rm s}(k)$ in Fig. (5a) and Fig. (5()b). This is indicative of the appearance of the strip-like intense vorticity patches in S2.
B.4 Pair dispersion
@@
We calculate how the separation, $R(t)$, between a pair of nearest-neighbor tracers, separated initially by the grid-spacing (approximately $0.3\eta$), evolves with time. We obtain $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations. Our results are shown in Fig. (8a). At early times, $\left\langle R^{2}(t)\right\rangle$ grows exponentially for all the runs [For S1 and S2 we show the plots in the inset of Fig. (8a)]. This range of exponential growth is very small for C1 and C2. Similar results [38] have been obtained for incompressible 2D turbulence. Exponential growth up to a length scale typically implies that the flow across the tracers is smooth up to that scale. This holds for our simulations because at such small time scales most of the pair separations lie away from shocks. We note that at late times the behavior of $\left\langle R^{2}\right\rangle$ is not universal.